# ZeroStereo: Zero-Shot Stereo Matching from Single Images

### v3

@@ Abstract
State-of-the-art supervised stereo matching methods have achieved remarkable performance on various benchmarks. However, their generalization to real-world scenarios remains challenging due to the scarcity of annotated real-world stereo data. In this paper, we propose ZeroStereo, a novel stereo image generation pipeline for zero-shot stereo matching. Our approach synthesizes high-quality right images from arbitrary single images by leveraging pseudo disparities generated by a monocular depth estimation model. Unlike previous methods that address occluded regions by filling missing areas with neighboring pixels or random backgrounds, we fine-tune a diffusion inpainting model to recover missing details while preserving semantic structure. Additionally, we propose Training-Free Confidence Generation, which mitigates the impact of unreliable pseudo labels without additional training, and Adaptive Disparity Selection, which ensures a diverse and realistic disparity distribution while preventing excessive occlusion and foreground distortion. Experiments demonstrate that models trained with our pipeline achieve state-of-the-art zero-shot generalization across multiple datasets with only a dataset volume comparable to Scene Flow. Code: https://github.com/Windsrain/ZeroStereo.
@@ 2 Related Work
Zero-shot Generalization in Stereo Matching. Despite these advancements, deep stereo models often struggle with generalization to real-world scenarios. DSMNet [62] addresses domain shifts by introducing domain normalization layers and non-local graph-based filters to enhance feature robustness. GraftNet [26] improves generalization by incorporating pre-trained features from large-scale datasets, while ITSA [10] mitigates shortcut learning using an information-theoretic approach. Inspired by masked representation learning, Rao et al. [38] propose a masking-based strategy to enhance stereo feature learning. Another line of research focuses on self-supervised learning using unlabeled images. Luo et al. [28] pioneers single-view stereo training on the KITTI dataset, while MfS-Stereo [53] generats stereo pairs from monocular images to enable training without ground-truth disparities. NeRF-Stereo [47] introduces NeRF to generate stereo images from 3D scene reconstructions.
@@ 3.3 Training-Free Confidence Generation
Assessing confidence in depth predictions remains challenging for previous monocular depth estimation models without additional training or auxiliary modules. Some post-processing methods rely on gradients [16] or probabilistic distributions [54] to estimate uncertainty.
Modern monocular depth estimation models [60, 61, 23] tend to predict relative depth, often represented as inverse depth in disparity space. This representation captures the relative distances between pixels without dependence on camera parameters. Therefore, when an image is flipped horizontally, the predicted relative depth between pixels is expected to remain unchanged.
@@ 3.4 Adaptive Disparity Selection
Figure 5: Visualization of different inpainting methods. Our fine-tuned SDv2I produces inpainted regions with minimal noise while preserving semantic structures that blend with the background.
@@ 3.5 Stereo Training
It is important to note that $\mathbf{I}_{l}^{r}$ includes pixels inpainted by the diffusion inpainting model. To exclude these pixels, we backward warp $1-\mathbf{M}_{inp}$ to obtain $\tilde{\mathbf{M}}_{inp}^{r}$. Besides, $\mathbf{I_{l}^{r}}$ may contain ghosting artifacts due to backward warping, which we filter using the non-occlusion mask $\mathbf{M}_{noc}$. Thus, our non-occlusion photometric loss is computed as follows:
@@ 4.1 Implementation Details
Diffusion Inpainting Model Fine-tuning. We utilize the Stable Diffusion V2 Inpainting (SDv2I) [39], disabling text conditioning and applying the DDPM noise scheduler [15] with 1,000 diffusion steps. Our experiments are conducted on a collection of synthetic stereo datasets, including Tartan Air [51], CREStereo Dataset [20], Scene Flow [29], VKITTI 2 [1], etc. Fine-tuning takes 50K steps with a batch size of 32 (gradient accumulation for 4 steps). We use the AdamW optimizer and a one-cycle learning rate schedule with a learning rate of 2e-5. Besides, we apply a crop size of $512\times 512$ and symmetric color augmentation.
Stereo Image Generation. We use the DDIM scheduler [43] and perform 50 sampling steps during inference. Following MfS-Stereo [53], we sample images from Depth in the Wild [4], COCO 2017 [24], DIODE [49], ADE20K [66], and Mapillary Vistas [34]. We randomly sample 35K of them to create a dataset called MfS35K. For disparity generation, we set the parameters as follows: $\mathbf{c}=0.1$, $\mathbf{r}=0.05$, $\mathbf{p}_{c}=0.8$, and $\mathbf{p}_{s}=\mathbf{p}_{l}=0.1$.
Stereo Matching Model Training. We utilize RAFT-Stereo [25] and IGEV-Stereo [56]. Models are trained on MfS35K with a batch size of 8 and a crop size of $384\times 512$. We follow all the source codes’ settings and train 200k steps from scratch. In addition to the data augmentation from RAFT-Stereo, we introduce Gaussian augmentation on the right image, as done in MfS-Stereo [53]. For ZeroStereo loss, we set $\beta=0.85,\mu=0.1$ same as NeRF-Stereo [47].
@@
RePaint [27] $256\times 256$ 2.7 156.5
StereoDiffusion [50] $512\times 512$ 14.6 31.2
Ours $512\times 512$ 5.8 1.9
@@ 4.2 Evaluation Datasets
KITTI 2015 [30] includes 200 training pairs with lidar ground truth for outdoor driving scenarios. ETH3D [41] contains 27 training pairs of grayscale images, covering outdoor and indoor scenarios. For Middlebury [40], we select the Middlebury V3 Benchmark Training Set (Midd-T), which consists of 15 training pairs for high-resolution indoor scenarios. During evaluation, we compute the endpoint error (EPE) and the percentage of pixels with errors greater than a threshold $\tau$.
4.3 Ablation Study
Figure 7: Visualization of different disparity selection.
Effectiveness of proposed modules. Tab.  1 shows the ablation study of our proposed modules. By adding ADS or TCG, we observe a notable reduction in EPE for both KITTI-15 and Midd-T. ADS expands the disparity selection range, improving the model’s ability to handle large disparities. As shown in Tab.  2, the disparity range adjusts adaptively according to the dataset resolution. TCG helps suppress unreliable labels, particularly in edges and textureless regions. However, adding Inpainting alone results in only a slight improvement. This finding supports our hypothesis that the fixed-range disparity selection in  [53] brings issues such as foreground distortion. As shown in Fig.  7, the large disparity ratio causes separation and distortion in the foreground which hinders the effectiveness of the diffusion inpainting model. When ADS is combined with Inpainting, we observe a significant performance improvement.
@@ 4.3 Ablation Study
Table 5: Comparision with NeRF-Stereo [47]. Models are trained with the same augmentation. Exception: ‡ official weights.
@@
When ADS, Inpainting, and TCG are all added, the performance consistently improves. Moreover, by introducing the final ZeroStereo Loss, the model can learn with $\mathcal{L}_{Zero}$ in low-confidence regions. This loss contributes to maintaining the model’s robustness, allowing it to achieve the best performance across different datasets. A more detailed analysis can be found in our supplementary materials.
Effectiveness of fine-tuned SDv2I. As shown in Fig.  5, pixels inpainted using our fine-tuned SDv2I preserve minimal noise and maintain the semantic structure closest to the background. As shown in Tab.  3, the fine-tuned SDv2I outperforms the pre-trained model. These results suggest that fine-tuning enhances the inpainting model’s ability to more accurately capture the semantic structure of the background.
Comparison with other synthesis methods. StereoDiffusion [50] introduces a training-free method for generating stereo images using the pre-trained SDv2. However, the inherent inconsistency between the stereo images arises because its warping operation is performed in the latent space. As shown in Fig.  6, StereoDiffusion suffers from structural distortions, texture inconsistencies, and poor occlusion handling, leading to unrealistic right-image generation. In addition, as shown in Tab.  4, our fine-tuned SDv2I is significantly faster and more memory-efficient than StereoDiffusion [50], while still achieving high-resolution synthesis. In contrast, RePaint [27] is tested at a lower resolution due to checkpoint limitations. A more detailed discussion is available in our supplementary materials.
@@ 4.4 Comparison with NeRF-Stereo
We compare our method with NeRF-Stereo [47], a leading method for generating stereo images. As shown in Tab.  5, all models are trained with our same data augmentation, except for NS-RAFT-Stereo33footnotemark: 3, which uses official weights. Since the official IGEV-Stereo limits $\mathcal{L}_{d}$ supervision to a disparity of 192, we split the Midd-T table by disparity to provide a more detailed evaluation.
Re-training NS-RAFT-Stereo with our data augmentation shows no improvement over the official weights, confirming that the gains achieved by our model are not solely due to augmentation. Zero-RAFT-Stereo outperforms NS-RAFT-Stereo33footnotemark: 3 by over 20% on average and better preserves background details, with only a 10% drop in Midd-T (F) for $\mathbf{D}<192$, whereas NS-RAFT-Stereo33footnotemark: 3 declines more, likely due to its dataset’s imperfect reconstruction of distant objects. Fig. 8 highlights NS-RAFT-Stereo’s failures in textureless regions, while our Zero-RAFT-Stereo shows over 40% improvement in handling such cases. Our model also excels in fine details, such as plants and holes. Moreover, as shown in Tab.  6, the model trained on MfS35K surpasses both the synthetic SceneFlow [29] and the NeRF-reconstructed NS65K [47], achieving the lowest edge error and demonstrating superior edge accuracy.
@@ 4.5 Zero-shot Generalization Benchmark
Following NeRF-Stereo [47], we construct a zero-shot generalization benchmark to compare our models with state-of-the-art methods. We include only those methods that have published official weights and evaluate them in the same environment. All methods are evaluated across the entire disparity range. For IGEV-Stereo [56], we train two versions under our pipeline: one using the original code settings for disparity supervision, and the other expanding the supervised range, consistent with RAFT-Stereo [25].
As shown in Tab.  7, our models demonstrate state-of-the-art zero-shot generalization performance across multiple datasets, both under the SceneFlow with ground truth (GT) and Real-world data without GT. Notably, Zero-RAFT-Stereo achieves the best or near-best results, particularly excelling in handling complex, real-world scenes.
@@ 5 Conclusion
We propose ZeroStereo, a novel stereo data generation pipeline for zero-shot stereo matching. The fine-tuned SDv2I adapts to complex inpainting masks and recovers background details. To handle unreliable pseudo labels, the TCG module leverages the spatial invariance of relative depth to compute confidence, helping to suppress uncertain labels. Besides, the ADS module generates a broader disparity distribution while avoiding foreground distortion. Finally, experimental results demonstrate that models trained under our pipeline achieve state-of-the-art zero-shot generalization performance.
@@
Acknowledgement. This research is supported by the National Key R&D Program of China (2024YFE0217700), National Natural Science Foundation of China (62472184, 623B2036), and the Fundamental Research Funds for the Central Universities.
@@ 7 Loss Analysis
In Sec.  3.5, we introduce the non-occlusion photometric loss $\mathcal{L}_{np}$ and the weighted final loss $\mathcal{L}_{Zero}$. However, their specific impact on stereo training has not been explicitly analyzed. As shown in Tab.  8, the methods listed from top to bottom correspond to: (1) applying the ordinary photometric loss $\mathcal{L}_{p}$, (2) using $\mathcal{L}_{p}$ with the weight $1-\mathbf{C}$, and (3) employing $\mathcal{L}_{np}$ with the weight $1-\mathbf{C}$.
@@ 8 Discussion on Synthesis Methods
StereoDiffusion [50] is a training-free method that utilizes a pre-trained latent diffusion model to generate stereo pairs from a single image. It applies null-text inversion [32] for image editing, first reversing the diffusion process to obtain a latent representation of the input image and then applying forward diffusion to synthesize the right view. However, this approach has notable limitations. First, inference is computationally expensive. As shown in Tab.  4, synthesizing a $512\times 512$ image takes approximately 30 seconds. Second, the null-text inversion process can unintentionally modify the left image, introducing content inconsistencies. As illustrated in Fig.  9, the original image lacks stones, yet both the generated left and right views erroneously include them. Similarly, fine details such as text often become distorted. Quantitative reconstruction loss measurements (Tab.  9) confirm these issues, showing significantly higher errors compared to our method. Moreover, using StereoDiffusion-generated stereo pairs for training stereo matching networks led to poor performance and convergence difficulties.
@@ 9 Additional Comparisons with NeRF-Stereo
For Midd-T, we report the performance of each sample in Tab.  11. Compared to NS-RAFT-Stereo [47], our Zero-RAFT-Stereo achieves improvements in nearly all cases. Notably, for samples where NS-RAFT-Stereo performs poorly, our method improves accuracy by almost 50%.
@@
Additionally, we evaluate both models on the DrivingStereo dataset under different weather conditions. As shown in Tab.  10, our Zero-RAFT-Stereo outperforms NS-RAFT-Stereo across all weather conditions except rainy weather, where both models exhibit poor performance, indicating a need for further optimization in such scenarios. Notably, Zero-RAFT-Stereo demonstrates significant improvements under foggy conditions, reducing errors from 18.18% to 8.66% at full resolution and from 3.41% to 1.70% at half resolution. Since foggy scenes typically have low contrast and poor visibility, these results suggest that Zero-RAFT-Stereo is more robust in such challenging conditions. As illustrated in Fig.  13, under extreme weather conditions, NS-RAFT-Stereo struggles to predict large textureless regions, while Zero-RAFT-Stereo successfully reconstructs complete ground surfaces and walls. Moreover, Zero-RAFT-Stereo exhibits superior segmentation of thin, tree-like objects and blurry background regions, highlighting its ability to maintain fine details even in adverse conditions.

### v4

@@ Abstract
State-of-the-art supervised stereo matching methods have achieved remarkable performance on various benchmarks. However, their generalization to real-world scenarios remains challenging due to the scarcity of annotated real-world stereo data. In this paper, we propose ZeroStereo, a novel stereo image generation pipeline for zero-shot stereo matching. Our approach synthesizes high-quality right images from arbitrary single images by leveraging pseudo disparities generated by a monocular depth estimation model. Unlike previous methods that address occluded regions by filling missing areas with neighboring pixels or random backgrounds, we fine-tune a diffusion inpainting model to recover missing details while preserving semantic structure. Additionally, we propose Training-Free Confidence Generation, which mitigates the impact of unreliable pseudo labels without additional training, and Adaptive Disparity Selection, which ensures a diverse and realistic disparity distribution while preventing excessive occlusion and foreground distortion. Experiments demonstrate that models trained with our pipeline achieve state-of-the-art zero-shot generalization across multiple datasets, with only a dataset volume comparable to Scene Flow. Code: https://github.com/Windsrain/ZeroStereo.
@@ 2 Related Work
Zero-shot Generalization in Stereo Matching. Despite these advancements, deep stereo models often struggle with generalization to real-world scenarios. DSMNet [64] addresses domain shifts by introducing domain normalization layers and non-local graph-based filters to enhance feature robustness. GraftNet [26] improves generalization by incorporating pre-trained features from large-scale datasets, while ITSA [10] mitigates shortcut learning using an information-theoretic approach. Inspired by masked representation learning, Rao et al. [38] propose a masking-based strategy to enhance stereo feature learning. Another line of research focuses on self-supervised learning using unlabeled images. Luo et al. [28] pioneers single-view stereo training on the KITTI dataset, while MfS-Stereo [54] generates stereo pairs from monocular images to enable training without ground-truth disparities. NeRF-Stereo [47] introduces NeRF to generate stereo images from 3D scene reconstructions.
@@ 3.3 Training-Free Confidence Generation
Assessing confidence in depth predictions remains challenging for previous monocular depth estimation models, which require additional training or auxiliary modules. Some post-processing methods rely on gradients [16] or probabilistic distributions [56] to estimate uncertainty.
Modern monocular depth estimation models [62, 63, 23] tend to predict relative depth, often represented as inverse depth in disparity space. This representation captures the relative distances between pixels, independent of camera parameters. Therefore, when an image is flipped horizontally, the predicted relative depth between pixels is expected to remain unchanged.
@@ 3.4 Adaptive Disparity Selection
Figure 5: Visualization of different inpainting methods.
@@ 3.5 Stereo Training
$\mathbf{I}_{l}^{r}$ includes pixels inpainted by the diffusion inpainting model. To exclude these pixels, we backward warp $1-\mathbf{M}_{inp}$ to obtain $\tilde{\mathbf{M}}_{inp}^{r}$. Besides, $\mathbf{I_{l}^{r}}$ contains ghosting artifacts due to backward warping, which we filter using $\mathbf{M}_{noc}$. Thus, our loss is computed as follows:
@@ 4.1 Implementation Details
Diffusion Inpainting Model Fine-tuning. We utilize the Stable Diffusion V2 Inpainting (SDv2I) [39], disabling text conditioning and applying the DDPM noise scheduler [15] with 1,000 diffusion steps. We use a collection of synthetic stereo datasets, including Tartan Air [52], CREStereo Dataset [20], Scene Flow [29], VKITTI 2 [1], etc. Fine-tuning takes 50K steps with a batch size of 32 (gradient accumulation for 4 steps). We use the AdamW optimizer and a one-cycle learning rate schedule with a learning rate of 2e-5. Besides, we apply a crop size of $\times 512$ 512 and symmetric color augmentation.
Stereo Image Generation. We use the DDIM scheduler [43] and perform 50 sampling steps. Following MfS-Stereo [54], we sample images from Depth in the Wild [4], COCO 2017 [24], DIODE [50], ADE20K [68], and Mapillary Vistas [34]. We randomly sample 35K to create a dataset called MfS35K. For disparity generation, we set: $\mathbf{c}=0.1$, $\mathbf{r}=0.05$, $\mathbf{p}_{c}=0.8$, and $\mathbf{p}_{s}=\mathbf{p}_{l}=0.1$.
Stereo Matching Model Training. We use RAFT-Stereo [25] and IGEV-Stereo [58]. Models are trained on MfS35K with a batch size of 8 and a crop size of 384×512384$\times$ 512384 × 512. We follow all the source codes’ settings and train 200k steps from scratch. In addition to the data augmentation from RAFT-Stereo, we introduce Gaussian augmentation on the right image, as done in MfS-Stereo [54]. For ZeroStereo loss, we set $\beta=0.85,\mu=0.1$ the same as NeRF-Stereo [47].
@@
RePaint [27] $\times 256$ 256 2.7 156.5
StereoDiffusion [51] $\times 512$ 512 14.6 31.2
Ours $\times 512$ 512 5.8 1.9
@@
Dataset Size KITTI-15 Midd-T ETH3D
>3px >2px >1px
Falling Things [49] 62K 5.14 5.81 28.63
CREStereo [20] 200K 6.26 10.91 2.56
Tartan Air [52] 307K 4.91 5.47 2.69
FoundationStereo [55] 1106K 4.69 5.11 2.52
MfS35K (Ours) 35K 4.53 4.45 2.13
Table 5: Ablation study of datasets.
Figure 7: Visualization of different disparity selection.
@@ 4.2 Evaluation Datasets
KITTI 2015 [30] includes 200 training pairs with lidar ground truth for outdoor driving scenarios. ETH3D [41] contains 27 training pairs of grayscale images, covering outdoor and indoor scenarios. For Middlebury [40], we select the Middlebury V3 Benchmark Training Set (Midd-T), which consists of 15 training pairs for high-resolution indoor scenarios.
@@
Table 6: Comparison with NeRF-Stereo [47]. Models are trained with the same augmentation. Exception: ‡ official weights.
@@
4.3 Ablation Study
Effectiveness of proposed modules. Tab.  1 shows the ablation study of our proposed modules. By adding ADS or TCG, we observe a notable reduction in EPE for both KITTI-15 and Midd-T. ADS improves the model’s ability to handle large disparities. As shown in Tab.  2, the disparity range adjusts adaptively according to the dataset resolution. TCG suppresses unreliable labels, particularly in edges and textureless regions. However, adding Inpainting alone results in only a slight improvement. As shown in Fig.  7, the large disparity ratio causes separation and distortion in the foreground, which hinders the effectiveness of the diffusion inpainting model. When ADS is combined with Inpainting, we observe a significant performance improvement.
When ADS, Inpainting, and TCG are all added, the performance consistently improves. Moreover, by introducing the final ZeroStereo Loss, the model can learn with $\mathcal{L}_{Zero}$ in low-confidence regions. This loss helps maintain the model’s robustness, enabling it to achieve optimal performance across various datasets. A more detailed analysis can be found in our supplementary materials.
Effectiveness of fine-tuned SDv2I. As shown in Fig.  5, pixels inpainted using our fine-tuned SDv2I preserve minimal noise and maintain the semantic structure closest to the background. As shown in Tab.  3, the fine-tuned SDv2I outperforms the pre-trained model. It suggests that fine-tuning enhances the inpainting model’s ability to capture the semantic structure of the background more accurately.
Comparison with other synthesis methods. StereoDiffusion [51] introduces a training-free method for generating stereo images using the pre-trained SDv2. However, the inherent inconsistency arises because the warping operation is performed in the latent space. As shown in Fig.  6, StereoDiffusion suffers from structural distortions, texture inconsistencies, and poor occlusion handling, leading to unrealistic right-image generation. As shown in Tab.  4, our fine-tuned SDv2I is significantly faster and more memory-efficient than StereoDiffusion [51], while still achieving high-resolution synthesis. A more detailed discussion is available in our supplementary materials.
Comparison with other synthetic datasets. In recent years, many synthetic datasets with better diversity and rendering realism have been proposed. As shown in Tab.  5, our MfS35K outperforms others, despite being significantly smaller. It reveals that rather than the absolute size of datasets, the diversity of scenarios is more beneficial for zero-shot generalization.
@@ 4.4 Comparison with NeRF-Stereo
We compare our method with NeRF-Stereo [47], a leading method for generating stereo images. As shown in Tab.  6, all models are trained with the same data augmentation, except for NS-RAFT-Stereo33footnotemark: 3, which uses official weights. Since the official IGEV-Stereo limits $\mathcal{L}_{d}$ supervision to a disparity of 192, we split the Midd-T table by disparity to provide a more detailed evaluation.
Re-training NS-RAFT-Stereo with our data augmentation shows no improvement, confirming that the gains are not solely due to augmentation. Zero-RAFT-Stereo outperforms NS-RAFT-Stereo33footnotemark: 3 by over 20%, with only a 10% drop in Midd-T (F) for $\mathbf{D}<192$, whereas NS-RAFT-Stereo33footnotemark: 3 declines more, likely due to its dataset’s imperfect reconstruction of distant objects. Fig. 8 highlights NS-RAFT-Stereo’s failures in textureless regions, while our Zero-RAFT-Stereo shows over 40% improvement in handling such cases. Moreover, as shown in Tab.  7, the model trained on MfS35K surpasses both the synthetic SceneFlow [29] and the NeRF-based NS65K [47], achieving the lowest edge error and superior edge accuracy.
@@ 4.5 Zero-shot Generalization Benchmark
Following NeRF-Stereo [47], we construct a zero-shot generalization benchmark. All methods are evaluated across the entire disparity range. For Zero-IGEV-Stereo, we train two versions: one using the original code settings for disparity supervision, and the other expanding the supervised range, consistent with RAFT-Stereo [25].
As shown in Tab.  8, our models demonstrate state-of-the-art zero-shot generalization performance across multiple datasets, both under the SceneFlow with ground truth (GT) and Real-world data without GT. Notably, Zero-RAFT-Stereo achieves the best or near-best results, particularly excelling in handling complex, real-world scenes.
@@ 5 Conclusion
We propose ZeroStereo, a novel stereo data generation pipeline for zero-shot stereo matching. The fine-tuned SDv2I adapts to complex inpainting masks and recovers background details. To handle unreliable pseudo labels, the TCG module leverages the spatial invariance of relative depth to compute confidence, helping to suppress uncertain labels. Besides, the ADS module generates a broader disparity distribution while avoiding foreground distortion. Finally, experiments demonstrate that our models achieve state-of-the-art zero-shot generalization performance.
@@
Acknowledgement. This research is supported by the National Key R&D Program of China (2024YFE0217700), National Natural Science Foundation of China (62472184, 623B2036), the Fundamental Research Funds for the Central Universities, and the Innovation Project of Optics Valley Laboratory (Grant No. OVL2025YZ005).
@@ 7 Loss Analysis
In Sec.  3.5, we introduce the non-occlusion photometric loss $\mathcal{L}_{np}$ and the weighted final loss $\mathcal{L}_{Zero}$. However, their specific impact on stereo training has not been explicitly analyzed. As shown in Tab.  9, the methods listed from top to bottom correspond to: (1) applying the ordinary photometric loss $\mathcal{L}_{p}$, (2) using $\mathcal{L}_{p}$ with the weight $1-\mathbf{C}$, and (3) employing $\mathcal{L}_{np}$ with the weight $1-\mathbf{C}$.
@@ 8 Discussion on Synthesis Methods
StereoDiffusion [51] is a training-free method that utilizes a pre-trained latent diffusion model to generate stereo pairs from a single image. It applies null-text inversion [32] for image editing, first reversing the diffusion process to obtain a latent representation of the input image and then applying forward diffusion to synthesize the right view. However, this approach has notable limitations. First, inference is computationally expensive. As shown in Tab.  4, synthesizing a $\times 512$ 512 image takes approximately 30 seconds. Second, the null-text inversion process can unintentionally modify the left image, introducing content inconsistencies. As illustrated in Fig.  9, the original image lacks stones, yet both the generated left and right views erroneously include them. Similarly, fine details such as text often become distorted. Quantitative reconstruction loss measurements (Tab.  10) confirm these issues, showing significantly higher errors compared to our method. Moreover, using StereoDiffusion-generated stereo pairs for training stereo matching networks led to poor performance and convergence difficulties.
@@ 9 Additional Comparisons with NeRF-Stereo
For Midd-T, we report the performance of each sample in Tab.  12. Compared to NS-RAFT-Stereo [47], our Zero-RAFT-Stereo achieves improvements in nearly all cases. Notably, for samples where NS-RAFT-Stereo performs poorly, our method improves accuracy by almost 50%.
@@
Additionally, we evaluate both models on the DrivingStereo dataset under different weather conditions. As shown in Tab.  11, our Zero-RAFT-Stereo outperforms NS-RAFT-Stereo across all weather conditions except rainy weather, where both models exhibit poor performance, indicating a need for further optimization in such scenarios. Notably, Zero-RAFT-Stereo demonstrates significant improvements under foggy conditions, reducing errors from 18.18% to 8.66% at full resolution and from 3.41% to 1.70% at half resolution. Since foggy scenes typically have low contrast and poor visibility, these results suggest that Zero-RAFT-Stereo is more robust in such challenging conditions. As illustrated in Fig.  13, under extreme weather conditions, NS-RAFT-Stereo struggles to predict large textureless regions, while Zero-RAFT-Stereo successfully reconstructs complete ground surfaces and walls. Moreover, Zero-RAFT-Stereo exhibits superior segmentation of thin, tree-like objects and blurry background regions, highlighting its ability to maintain fine details even in adverse conditions.